Need help solving this problem
Solve for x: (The following is all the info I have about this right triangle.)
Leg a is x
Leg b is 4x
Leg c is ?

Answer by stanbon(75887)

Solve for x: (The following is all the info I have about this right triangle.)
Leg a is x
Leg b is 4x
Leg c is ?
---------------------
The answer depends on whether leg c is or is not the hypotenuse.
-----------
1st If it is the hypotenuse:
c^2 = a^2 + b^2
c^2 = x^2 + (4x)^2 = 17x^2
c = xsqrt(17)
=======================
2nd: if it is not the hypotenuse and b is the hypotenuse:
b^2 = a^2 + c^2
(4x)^2 = x^2 + c^2
c^2 = 16x^2 - x^2
c^2 = 15x^2
c = xsqrt(15)
=========================
Cheers,
Stan H.